OpenStudy (anonymous):

Use half angle identity to find the exact value of sin 112.5º?

OpenStudy (yanasidlinskiy):

\(\Huge\cal \color{turquoise}{Welcome~\to~OpenStudy!!}\hspace{-310pt}\color{yellow}{Welcome~\to~OpenStudy!!}\hspace{-307.1pt}\color{#ff7100}{Welcome~\to~\color{#00B4ff}{Open}}\color{#7cc517}{Study!!}\)

zepdrix (zepdrix):

Ooo that's colorful ^^ So we can start with this: \[\Large\rm \sin\left(112.5\right)=\sin\left(\frac{225}{2}\right)\]

OpenStudy (anonymous):

But how did you get to sin(225/2)?

zepdrix (zepdrix):

So then we need to use our Half-Angle Identity for Sine,\[\Large\rm \sin\left(\frac{x}{2}\right)=\pm\sqrt{\frac{1-\cos(x)}{2}}\]

OpenStudy (yanasidlinskiy):

Yea...I know..I was going to help but since you already startrd...I'm going to let you finish!

zepdrix (zepdrix):

Ummmm

zepdrix (zepdrix):

I multiplied \(\Large\rm 112.5\) by \(\Large\rm \dfrac{2}{2}\) The reason I did this is because we really need a 2 denominator to apply our Half-Angle Formula.

zepdrix (zepdrix):

\[\Large\rm 112.5\cdot\frac{2}{2}=\frac{112.5}{1}\cdot \frac{2}{2}=\frac{2\cdot112.5}{2}=\frac{225}{2}\]Ok with that part? :o We got that 2 denominator we were looking for.

OpenStudy (anonymous):

Yes, I get that but then you just substitute for your half angle identity?

zepdrix (zepdrix):

\[\Large\rm \sin\left(\frac{\color{orangered}{225}}{2}\right)=\pm\sqrt{\frac{1-\cos(\color{orangered}{225})}{2}}\]Yes :) Is 225 one of our special angles? Blah I can't remember. If it's not, we'll have to apply Angle Sum/Difference Formula.

zepdrix (zepdrix):

Oh it is, ok good :)

zepdrix (zepdrix):

You don't want to leave the plus/minus sign in front though! Look back at your original angle, and determine which quadrant it is in. And then try to remember if Sine is positive or negative in that quadrant.

zepdrix (zepdrix):

Example: If we had started with \(\Large\rm \sin(193.4^{\circ})\) We would only want the `negative` root, because sine is negative in the third quadrant.

OpenStudy (anonymous):

The original angle is 112.5, right? which would be in the second quadrant, so it would be positive?

zepdrix (zepdrix):

Mmm yes very good. So we only want the positive root. We can drop the plus/minus sign.\[\Large\rm \sin\left(\frac{\color{orangered}{225}}{2}\right)=\sqrt{\frac{1-\cos(\color{orangered}{225})}{2}}\]

OpenStudy (anonymous):

Going back to 225º though, what do you mean by it being a special angle?

zepdrix (zepdrix):

Remember that at special places around the unit circle sine and cosine are giving us values like:\[\Large\rm \frac{1}{2}\]\[\Large\rm \frac{\sqrt2}{2}\]\[\Large\rm \frac{\sqrt3}{2}\]Those look familiar? :o

OpenStudy (anonymous):

sine of 30, 45 and 60

zepdrix (zepdrix):

Mmm yah :) Those values continue around the circle. So ummm.. I guess one way you can do it is....\[\Large\rm \cos(225)=-\cos(225-180)=-\cos(45)\]So what I did here is......

OpenStudy (anonymous):

you used reference angle

zepdrix (zepdrix):



zepdrix (zepdrix):

Yah if you remember your reference angles you can do a lot of stuff :p

zepdrix (zepdrix):

If this method is too confusing, I apologize >.< So what I was doing was, spinning halfway around the circle, putting a negative in front since our sign on cosine changed (from Q3 to Q1) and relating it to 45 degrees.

OpenStudy (anonymous):

so then cos 45º is sqrt of 3/2?

zepdrix (zepdrix):

Nooo you silly billy. You really gotta try to remember the first quadrant. 45 degrees is pi/4. I find it easier to remember the angles in radians myself. All of the pi/4 thingies give us sqrt2/2, For both sine and cosine. just have to watch the `sign` on it.

OpenStudy (anonymous):

oh yes, its true, i'm sorry!

OpenStudy (anonymous):

I think i made a mess though. Can we go back to our half angle identity?

zepdrix (zepdrix):

Are we ok up to this point?\[\Large\rm \sin\left(\frac{\color{orangered}{225}}{2}\right)=\sqrt{\frac{1-\cos(\color{orangered}{225})}{2}}\]Or we need to back up a lil further?

OpenStudy (anonymous):

thats okay, continue please

OpenStudy (loser66):

support @zepdrix 112.5= 5pi/8

zepdrix (zepdrix):

In radians? :) yah that's way to do it also.

zepdrix (zepdrix):

So we're trying to figure out cos(225). Having trouble there?

OpenStudy (anonymous):

no, 225-180, ends up as 45º

zepdrix (zepdrix):

so what do we get for cos(225)? :)\[\Large\rm \cos(225)=?\]

OpenStudy (anonymous):

-sqrt2/2

zepdrix (zepdrix):

Ok great! Let's plug that in. \[\Large\rm \sin\left(\frac{\color{orangered}{225}}{2}\right)=\sqrt{\frac{1-(-\sqrt2/2)}{2}}\]